# An operator learning perspective on parameter-to-observable maps

### v1

@@ 4.3.1. Discussion
The proof of Theorem 4.7 is based on a bias–variance decomposition argument that is detailed in Appendix C.2. Notice that the error bound in the theorem is uniform over $\mathcal{H}^{s}$–balls because the implied constant in the inequality (29) does not depend on $f^{\dagger}$. The probability that the bound fails to hold decays to zero faster than any power law as a function of the sample size $N$ because $\alpha+s-1>0$ by hypothesis. The convergence rate (29) depends on the three smoothness exponents $\alpha$, $\alpha^{\prime}$, and $s$. The consequences are the same as those identified in [16, Section 3] because the error bound takes the same form. Indeed, rougher training data points $\{u_{n}\}_{n=1}^{N}$ (smaller $\alpha$), smoother test data points $u^{\prime}\sim\nu^{\prime}$ (larger $\alpha^{\prime}$), and smoother target functionals $f^{\dagger}$ (larger $s$) all serve to reduce the test error (up to saturation).
@@
Theorem A.1 bounds the conditional expectation of the first term on the right hand side of the preceding equality. We see that the only new error term that the full posterior introduces is the second term, the posterior spread. But the end of Subsection C.2.1 explains that the posterior spread may be upper bounded by a constant times the rate $\varepsilon_{N}^{2}$ from Theorem A.1. Thus, the end-to-end error bounds (54) and (29) remain valid for the posterior sample estimator $f^{(N)}$ at the expense of enlarged constant factors. Posterior contraction rates then follow from a standard Chebyshev inequality argument [16, Section 3.3, p. 18]. Similar results may be deduced for the full-field setting (FF) because the error analysis for the forward map in [16, Section 3.4] already takes into account the full posterior distribution (27).
@@ 4.3.2. Sample complexity comparison
To conclude Subsection 4.3, we provide a detailed comparison of the end-to-end (EE) and full-field (FF) PtO map learning approaches to provide intuition. Focusing on the specific setting of QoIs with power law coefficient decay and in-distribution test error, the following corollary is a consequence of Theorem 4.7 and Theorem 4.8. A short proof is provided in Appendix C.4.
@@
Instate the notation and assertions in Assumptions 4.2, 4.3, and (A-III). Suppose that the training and test distributions are equivalent, i.e., $\alpha^{\prime}=\alpha$. Let the underlying true PtO map $f^{\dagger}$ have the factorization $f^{\dagger}=q^{\dagger}\circ L^{\dagger}$, where $\lvert q^{\dagger}(\varphi_{j})\rvert^{2}\lesssim j^{-2r-1}$ as $j\to\infty$ and $L^{\dagger}$ is as in (26) with eigenvalues $l^{\dagger}\in\mathcal{H}^{\beta}$. If $\beta+r+1/2>0$, $\alpha+\beta+r>1/2$, and $\alpha+\beta>0$, then there exist constants $c>0$ and $C>0$ such that for all sufficiently large $N$, the following holds on an event with probability at least $1-C\exp(-cN^{\min(1,\frac{\alpha+\beta+r-1/2}{1+\alpha+\beta+r})})$ over $U=\{u_{n}\}_{n=1}^{N}\sim\nu^{\otimes N}$. The (EE) posterior mean estimator $\bar{f}^{(N)}$ in (15) (with $p\coloneqq\beta+r+1$ in (A3)) trained on end-to-end data $(U,Y)$ satisfies
@@
Several interesting insights may be deduced from the convergence rates (34) and (35). Since a common set of assumptions have been identified in the statement of Corollary 4.10, these rates may be directly compared to assess whether the (EE) or (FF) approach is more accurate or, equivalently, more data efficient, than the other. First, we note that the squared generalization error of (EE) in (34) has a nonparametric convergence rate $N^{-(1-\delta)}$ that is always slower than the parametric estimation rate $N^{-1}$ by a polynomial factor $N^{\delta}$ (where $\delta>0$). On the other hand, if $r\geq 0$, then (FF) achieves the fast parametric rate $N^{-1}$ (up to a log factor if $r=0$); this always beats (EE) in the regime $r\geq 0$. This regime has an interesting interpretation because the QoI is continuous if $r>0$. These types of QoIs appear naturally in scientific applications.
@@
By the concavity of $r\mapsto\rho_{\mathrm{EE}}(r)$ in the range $[r_{0},r_{1}]$ and affine structure of $r\mapsto\rho_{\mathrm{FF}}(r)$, we deduce the following two insights:
@@
The inequalities in (I1) and (I2) suggest that the (FF) approach is advantageous when the QoI is smooth and the (EE) approach is advantageous when the QoI is rough. An example of a rough QoI is $q^{\dagger}\colon h\mapsto(dh/dx)(x_{0})$, which returns a point evaluation of the first derivative of a univariate function (see the last item in Remark 4.6). Direct application of such a rough QoI to a function is an ill-posed operation (e.g., amplifies perturbations in the function). This may partially explain why (EE) is preferable in this case, as the (EE) estimator does not require the evaluation of $q^{\dagger}$ while (FF) does. We plot the functions $\rho_{\mathrm{EE}}$ and $\rho_{\mathrm{FF}}$ in Figure 2, which visualizes the main insights (I1) and (I2) from the preceding discussion.
@@
Figure 2. (EE) vs. (FF) convergence rate exponents (36) as a function of QoI decay exponent $r$. Larger exponents imply faster convergence rates. As the curves gets lighter, $\alpha+\beta$, an indicator of the smoothness of the problem, increases. The vertical dashed line corresponds to $r=-1/2$, which is the transition point where (EE) and (FF) have the same rate and the onset of power law decay for the QoI coefficients begins.
@@ 5. Numerical experiments
We now perform numerical experiments with the proposed FNM architectures. These experiments have two main purposes. The first is to numerically implement and compare the various FNM models on several PtO maps of practical interest; the second is to qualitatively validate the theory developed in the paper. We focus on nontrivial nonlinear problems with finite-dimensional observables that define the QoI maps. Although our linear theory from Section 4 does not apply to such nonlinear problems, we still observe qualitative validation of the main implications of the linear analysis. That is, for smooth enough QoIs, full-field learning is at least as data-efficient as end-to-end learning. Unlike the theory, however, our numerical results distinguish the two approaches only by constant factors and not by the actual convergence rates.
@@
over the FNM’s tunable parameters $\theta$ using mini-batch SGD with the ADAM optimizer. The choice of the loss function is dependent on the underlying problem. Moreover, the norm in the preceding displays are inferred from the space that the $\tilde{y}_{n}$ takes values in (i.e., finite-dimensional or infinite-dimensional output spaces). To avoid numerical instability in our actual computations, we add $10^{-6}$ to the denominator of the ratio in (37).
The numerical experiments are organized as follows. In Subsection 5.1, we extract the first four polynomial moments from the solution of a velocity-parametrized 2D advection-diffusion equation. Next, Subsection 5.2 considers the flow over an airfoil modeled by the steady compressible Euler equation. The PtO map sends the shape of the airfoil to the resultant drag and lift force vector. Last, we study an elliptic homogenization problem parametrized by material microstructure in Subsection 5.3. Here, the QoI returns the effective tensor of the material.

### v2

@@ 4.3.1. Discussion
The proof of Theorem 4.8 is based on a bias–variance decomposition argument that is detailed in Appendix C.2. Notice that the error bound in the theorem is uniform over $\mathcal{H}^{s}$-balls because the implied constant in the inequality (33) does not depend on $f^{\dagger}$. The probability that the bound fails to hold decays to zero faster than any power law as a function of the sample size $N$ because $\alpha+s-1>0$ by hypothesis. The convergence rate (33) depends on the three smoothness exponents $\alpha$, $\alpha^{\prime}$, and $s$. The consequences are the same as those identified in [17, Section 3] because the error bound takes the same form. Indeed, rougher training data points $\{u_{n}\}_{n=1}^{N}$ (smaller $\alpha$), smoother test data points $u^{\prime}\sim\nu^{\prime}$ (larger $\alpha^{\prime}$), and smoother target functionals $f^{\dagger}$ (larger $s$) all serve to reduce the test error (up to saturation).
@@
Theorem A.1 bounds the conditional expectation of the first term on the right-hand side of the preceding equality. We see that the only new error term that the full posterior introduces is the second term, the posterior spread. But the end of Subsection C.2.1 explains that the posterior spread may be upper bounded by a constant times the rate $\varepsilon_{N}^{2}$ from Theorem A.1. Thus, the end-to-end error bounds (59) and (33) remain valid for the posterior sample estimator $f^{(N)}$ at the expense of enlarged constant factors. Posterior contraction rates then follow from a standard Chebyshev inequality argument [17, Section 3.3, p. 18]. Similar results may be deduced for the full-field setting (FF) because the error analysis for the forward map in [17, Section 3.4] already takes into account the full posterior distribution (27).
@@ 4.3.2. Sample complexity comparison
To conclude Subsection 4.3, we provide a detailed comparison of the end-to-end (EE) and full-field (FF) PtO map learning approaches to provide intuition about their statistical performance. Focusing on the specific setting of QoIs with power law coefficient decay and in-distribution test error (i.e., $\alpha^{\prime}=\alpha$), the following corollary is a consequence of Theorem 4.8 and Theorem 4.9. A short proof is provided in Appendix C.4.
@@
Instate the notation and assertions in Assumptions 4.2, 4.3, and (A-III). Suppose that the training and test distributions are equal, i.e., $\nu=\nu^{\prime}$. Let the underlying true PtO map $f^{\dagger}$ have the factorization $f^{\dagger}=q^{\dagger}\circ L^{\dagger}$, where $\lvert q^{\dagger}(\varphi_{j})\rvert^{2}\lesssim j^{-2r-1}$ as $j\to\infty$ and $L^{\dagger}$ is as in (26) with eigenvalues $l^{\dagger}\in\mathcal{H}^{\beta}$. If $\beta+r+1/2>0$, $\alpha+\beta+r>1/2$, and $\alpha+\beta>0$, then there exist constants $c>0$ and $C>0$ such that for all sufficiently large $N$, the following holds on an event with probability at least $1-C\exp(-cN^{\min(1,\frac{\alpha+\beta+r-1/2}{1+\alpha+\beta+r})})$ over $U=\{u_{n}\}_{n=1}^{N}\sim\nu^{\otimes N}$. The (EE) posterior mean estimator $\bar{f}^{(N)}$ in (15) (with $p\coloneqq\beta+r+1$ in (A3)) trained on end-to-end data $(U,Y)$ satisfies
@@
Several interesting insights may be deduced from the convergence rates (38) and (39). Since a common set of assumptions have been identified in the statement of Corollary 4.11, these rates may be directly compared to assess whether the (EE) or (FF) approach is in general more accurate or, equivalently, more data-efficient, than the other—at least up to the sharpness of the upper bounds. First, we note that the squared generalization error of (EE) in (38) has a nonparametric convergence rate $N^{-(1-\delta)}$ that is always slower than the parametric estimation rate $N^{-1}$ by a polynomial factor $N^{\delta}$ (where $\delta>0$). On the other hand, if $r\geq 0$, then (FF) achieves the fast parametric rate $N^{-1}$ (up to a log factor if $r=0$); this always beats the (EE) upper bound in the regime $r\geq 0$. This regime has an interesting interpretation because the QoI is continuous if $r>0$. These types of QoIs appear naturally in scientific applications.
@@
By the concavity of $r\mapsto\rho_{\mathrm{EE}}(r)$ in the range $[r_{0},r_{1}]$ and affine structure of $r\mapsto\rho_{\mathrm{FF}}(r)$, we deduce the following two insights from the standpoint of upper bounds:
@@
Figure 2. (EE) vs. (FF) convergence rate exponents (40) as a function of QoI regularity exponent $r$. Larger exponents imply faster convergence rates. As the curves gets lighter, $\alpha+\beta$, an indicator of the smoothness of the problem, increases. The vertical dashed line corresponds to $r=-1/2$, which is the transition point where (EE) and (FF) have the same rate and the onset of power law decay for the QoI coefficients begins.
The inequalities in (I1) and (I2) suggest that the (FF) approach is advantageous when the QoI is smooth and the (EE) approach is advantageous when the QoI is rough. Numerical evidence in the next subsection strongly supports this assertion. However, definitive claims would require lower bounds to precisely separate the two methods. An example of a rough QoI is $q^{\dagger}\colon h\mapsto(dh/dx)(x_{0})$, which returns a point evaluation of the first derivative of a univariate function; see the last item in Remark 4.7. Direct application of such a rough QoI to a function is an ill-posed operation, e.g., amplifies perturbations in the function. This may partially explain why (EE) is preferable in this case, as the (EE) estimator does not require the evaluation of $q^{\dagger}$ while (FF) does. We plot the functions $\rho_{\mathrm{EE}}$ and $\rho_{\mathrm{FF}}$ in Figure 2, which visualizes the main insights (I1) and (I2) from the preceding discussion.
4.4. Simulation study
To validate the theoretical results in Subsection 4.3, we numerically implement the linear (EE) and (FF) estimators on a synthetic data benchmark problem. Specializing to the setting of Corollary 4.11, we take $H=L^{2}((0,1);\mathbb{R})$ with orthonormal basis $\{\varphi_{j}\}_{j\in\mathbb{N}}$ defined by $\varphi_{j}(x)=\sqrt{2}\sin(j\pi x)$ for each $x\in(0,1)$ and $j\in\mathbb{N}$. The forward map $L^{\dagger}$ is defined by its diagonalization (26) with $l^{\dagger}_{j}\coloneqq(j\pi)^{-2}$ for each $j\in\mathbb{N}$. In physical space, this corresponds to the solution operator of the one-dimensional Poisson differential equation equipped with Dirichlet boundary conditions. It follows that $l^{\dagger}\in\mathcal{H}^{\beta}(\mathbb{N};\mathbb{R})$ for any $\beta<\beta^{\star}\coloneqq 3/2$. We fix $L^{\dagger}$ in what follows. Next, we vary the ground truth QoI map $q^{\dagger}$ across the three examples in Remark 4.7—a point evaluation of a derivative, a point evaluation, and an integral. These QoIs satisfy the decay condition $\lvert q^{\dagger}(\varphi_{j})\rvert^{2}\lesssim j^{-2r-1}$ for $r=-3/2$, $r=-1/2$, and $r=1/2$, respectively. The resulting true factorized linear functional $f^{\dagger}\coloneqq q^{\dagger}\circ L^{\dagger}$ has coordinates $f^{\dagger}_{j}\coloneqq f^{\dagger}(\varphi_{j})=q^{\dagger}(\varphi_{j})l^{\dagger}_{j}$ for each $j\in\mathbb{N}$.
The data and prior covariance operators are diagonalized in the basis $\{\varphi_{j}\}_{j\in\mathbb{N}}$ and are characterized by their eigenvalues, which for each $j\in\mathbb{N}$ belong to the set
${\mathopen{}\mathclose{{}\left\{\tau^{2\vartheta-1}(\pi^{2}j^{2}+\tau^{2})^{-\vartheta}\,\middle|\,\tau\in\mathbb{R}_{>0}\quad\text{and}\quad\vartheta\in\mathbb{R}}\right\}}\,.$ (41)
This spectrum corresponds to a Matérn-like covariance structure. For the eigenvalues $\{\sigma_{j}\}_{j\in\mathbb{N}}$ of the data covariance operator $\Sigma$ (11), we fix $\tau=15$ and $\vartheta=0.75$ in (41). This implies that the asymptotic data decay assumption (A2) is satisfied with $\alpha=\vartheta=0.75$; the numerical results are qualitatively the same when $\alpha$ equals $2$ or $4.5$. The hypotheses $\beta+r+1/2>0$, $\alpha+\beta+r>1/2$, and $\alpha+\beta>0$ of Corollary 4.11 are always satisfied given the above values of $\alpha=0.75$, $\beta<3/2$, and $r\in\{-3/2,-1/2,1/2\}$. Next, we choose the data distribution $\nu$ to be a Gaussian measure $\mathcal{N}(0,\Sigma)$, which has a KL expansion (22) as in Assumption 4.3 with i.i.d. coefficients $z_{j}\sim\mathcal{N}(0,1)$ for each $j$. For the Gaussian prior distributions, in (41) we fix $\tau=2$ and take $\vartheta=\beta+r+1$ for the (EE) estimator prior (19) and $\vartheta=\beta+1/2$ for the (FF) estimator prior (A-III), as required by the hypotheses of Corollary 4.11. Finally, we take the scalar standard deviation parameter for the Gaussian white noise processes that corrupt the observed outputs to be $\gamma=10^{-3}$ for both the (EE) and (FF) training datasets.
Recalling the prior covariance operator $\Lambda$ (19), the input training data $\{u_{n}\}_{n=1}^{N}\sim\nu^{\otimes N}$, and the sampling operator $S_{N}$ from Subsection 4.1.1, we efficiently implement the (EE) posterior mean estimator (15) in dual form via the explicit formula
$\bar{f}^{(N)}=\Lambda S_{N}^{*}\bigl{(}S_{N}^{\phantom{*}}\Lambda S_{N}^{*}+\gamma^{2}\operatorname{Id}_{\mathbb{R}^{N}}\bigr{)}^{-1}Y\,.$ (42)
The preceding display follows from (15), (16), and an infinite-dimensional Woodbury matrix identity [65, Theorem 1]. Working in coordinates, we obtain for each $j\in\mathbb{N}$ the convenient sequence space representation
$\bar{f}^{(N)}_{j}\coloneqq\langle\bar{f}^{(N)},\varphi_{j}\rangle_{L^{2}}=\lambda_{j}\sum_{n=1}^{N}u_{jn}V_{n}^{(N)}\,,$ (43)
where $\{\lambda_{j}\}_{j\in\mathbb{N}}$ are the eigenvalues of $\Lambda$, $u_{jn}\coloneqq\langle u_{n},\varphi_{j}\rangle_{L^{2}}$, and $V^{(N)}\coloneqq(S_{N}^{\phantom{*}}\Lambda S_{N}^{*}+\gamma^{2}\operatorname{Id}_{\mathbb{R}^{N}})^{-1}Y\in\mathbb{R}^{N}$. The Gram matrix $S_{N}^{\phantom{*}}\Lambda S_{N}^{*}\in\mathbb{R}^{N\times N}$ has entries $\sum_{j\in\mathbb{N}}\lambda_{j}u_{jn}u_{jn^{\prime}}$ for each $n\in[N]$ and $n^{\prime}\in[N]$. We discretize these entries by spectral truncation—which replaces sums over the countable set $\mathbb{N}$ by sums over the finite set $\{j\in\mathbb{N}\colon j\leq J\}$—and we also retain only the first $J$ modes in (43) for fixed $J=4096$. The (FF) estimator $\tilde{f}^{(N)}\coloneqq q^{\dagger}\circ\bar{L}^{(N)}$ (28) is defined by its coordinates $\tilde{f}^{(N)}_{j}\coloneqq\tilde{f}^{(N)}(\varphi_{j})=q^{\dagger}(\varphi_{j})\bar{l}_{j}^{(N)}$, where $\bar{l}_{j}^{(N)}$ is as in (27). We also discretize this estimator by spectral truncation with $J=4096$ resolved Fourier modes. The results for both the (EE) and (FF) estimators are nearly identical if the resolution $J$ is increased to $8192$ or higher.
(a) $r=-3/2$
(b) $r=-1/2$
(c) $r=1/2$
Figure 3. Empirical sample complexity of the Bayesian (EE) and (FF) estimators for linear PtO maps based on a Poisson problem. The solid purple lines are best linear fits to the broken curves with markers, which correspond to numerically computed squared errors. In all three figures, the experimentally observed convergence rates are nearly perfect matches to those from the theoretical upper bounds in Corollary 4.11 (see Table 1).
In Figure 3, we plot the expected relative squared $L^{2}_{\nu}(H;\mathbb{R})$ Bochner norm error as a function of the sample size $N$. Specifically, the vertical axis represents
$\frac{\operatorname{\mathbb{E}}\Bigl{[}\operatorname{\mathbb{E}}^{u\sim\nu}\lvert\langle f^{\dagger},u\rangle_{L^{2}}-\langle\widehat{f},u\rangle_{L^{2}}\rvert^{2}\Bigr{]}}{\operatorname{\mathbb{E}}^{u\sim\nu}\lvert\langle f^{\dagger},u\rangle_{L^{2}}\rvert^{2}}=\frac{\operatorname{\mathbb{E}}\Bigl{[}\sum_{j=1}^{\infty}\sigma_{j}\lvert f^{\dagger}_{j}-\widehat{f}_{j}\rvert^{2}\Bigr{]}}{\sum_{k=1}^{\infty}\sigma_{k}\lvert f^{\dagger}_{k}\rvert^{2}}\,,$ (44)
where $\widehat{f}$ is either the (EE) or (FF) estimator, $\widehat{f}_{j}\coloneqq\langle\widehat{f},\varphi_{j}\rangle_{L^{2}}$, and the outer expectation is with respect to the realizations of the $N$ training data pairs. For each $N\in\{2^{4},\ldots,2^{14}\}$, we approximate the outer expectation in the preceding display by averaging over $1000$ i.i.d. realizations of the training data; hence, $1000$ estimators are trained for each $N$. The error bounds in Corollary 4.11 remain valid for the full outer expectation in (44) by [62, Corollary D.2, p. 226]. The infinite series in (44) are discretized by finite sums over the first $J=4096$ summands. The solid purple lines in Figure 3 are linear least squares fits to the logarithm of the computed squared errors. We then take the slopes of these lines as the experimental convergence rate exponents, which are recorded in parentheses in Table 1.
Table 1. Sample complexity comparison experiment. The entries of the table record the theoretical vs. experimental (in parentheses) convergence rate exponents $\rho$ in $O(N^{-\rho})$ of the relative expected squared $L_{\nu}^{2}(H;\mathbb{R})$ in-distribution error (44). The theoretical rates are taken from the upper bound in Corollary 4.11. Here $\alpha=0.75$.
$r=-3/2$ $r=-1/2$ $r=1/2$
$\mathrm{EE}$ 0.714 (0.738) 0.818 (0.815) 0.867 (0.864)
$\mathrm{FF}$ 0.455 (0.455) 0.818 (0.818) 1.000 (1.003)
Table 1 strongly suggests that for the specific linear functional learning problem under consideration, the convergence rates from the upper bounds in Corollary 4.11 are sharp. Indeed, the empirical numerical results are in near perfect agreement with our theoretical predictions for all three QoI maps of varying smoothness. In particular, the (EE) estimator indeed converges faster than the (FF) one for the rough derivative point evaluation QoI (32). For the point evaluation QoI (31), the rates are identical, as predicted. The (FF) estimator converges at the fast $N^{-1}$ rate for the mean QoI (30), which is smooth, while the (EE) estimator converges strictly slower. Going beyond the asymptotic theory, Figure 3 also shows that for fixed $N$, the estimator with a faster convergence rate also has a smaller expected relative squared error in the regime of sample sizes considered, i.e., better constants.
@@ 5. Numerical experiments
We now perform numerical experiments with the proposed FNM architectures. These experiments have two main purposes. The first is to numerically implement and compare the various FNM models on several PtO maps of practical interest; the second is to qualitatively validate the theory developed in the paper for such maps. We focus on nontrivial nonlinear problems with finite-dimensional observables that define the QoI maps. Although our linear theory from Section 4 does not apply to such nonlinear problems, we still observe qualitative validation of the main implications of the linear analysis. That is, for smooth enough QoIs, full-field learning is at least as data-efficient as end-to-end learning. Unlike the theory, however, our numerical results distinguish the two approaches only by constant factors and not by the actual convergence rates.
@@
over the FNM’s tunable parameters $\theta$ using mini-batch SGD with the ADAM optimizer. The choice of the loss function is dependent on the underlying problem. Moreover, the norm in the preceding displays are inferred from the space that the $\tilde{y}_{n}$ takes values in (i.e., finite-dimensional vector or infinite-dimensional function output spaces). To avoid numerical instability in our actual computations, we add $10^{-6}$ to the denominator of the ratio in (45).
Remark 5.1 (data discretization error).
In addition to the discretization error introduced by the discrete and implementable realizations of the continuum FNM architectures [44], there is another source of discretization error due to our choice of data generation procedure. Specifically, the training and test data in this section are generated from numerical solvers that discretely approximate an underlying continuum operator at a fixed resolution. The weights of the resulting trained FNMs have a complicated dependence on this discretization error. Although simpler operator learning architectures are stable to such errors [42, Example 3.9, pp. 5–6], no such results exist yet for neural operators. Furthermore, in line with most of the literature, the empirical convergence results that we numerically report in this section are for the test error with respect to the discretized operator or PtO map. Thus, there is an implicit assumption that this discretized operator is sufficiently resolved so that the computable but discrete test error is an accurate surrogate for the true but inaccessible test error with respect to the continuum operator. Alternative data acquisition strategies may mitigate these effects to some extent [26].
The numerical experiments are organized as follows. In Subsection 5.1, we extract the first four polynomial moments from the solution of a velocity-parametrized 2D advection–diffusion equation. Next, Subsection 5.2 considers the flow over an airfoil modeled by the steady compressible Euler equation. The PtO map sends the shape of the airfoil to the resultant drag and lift force vector. Last, we study an elliptic homogenization problem parametrized by material microstructure in Subsection 5.3. Here, the QoI returns the effective tensor of the material.